# NoPain: No-box Point Cloud Attack via Optimal Transport Singular Boundary

### v2

@@ Abstract
Adversarial attacks exploit the vulnerability of deep models against adversarial samples. Existing point cloud attackers are tailored to specific models, iteratively optimizing perturbations based on gradients in either a white-box or black-box setting. Despite their promising attack performance, they often struggle to produce transferable adversarial samples due to overfitting the specific parameters of surrogate models. To overcome this issue, we shift our focus to the data distribution itself and introduce a novel approach named NoPain, which employs optimal transport (OT) to identify the inherent singular boundaries of the data manifold for cross-network point cloud attacks. Specifically, we first calculate the OT mapping from noise to the target feature space, then identify singular boundaries by locating non-differentiable positions. Finally, we sample along singular boundaries to generate adversarial point clouds. Once the singular boundaries are determined, NoPain can efficiently produce adversarial samples without the need of iterative updates or guidance from the surrogate classifiers. Extensive experiments demonstrate that the proposed end-to-end method outperforms baseline approaches in terms of both transferability and efficiency, while also maintaining notable advantages even against defense strategies. The source code will be publicly available.
@@ 2 Related Work
White-box attack on point cloud. Existing works of point cloud attacks can be roughly divided into white-box attacks and black-box attacks. White box attack means that all information of the attacked classification model can be used to update perturbations, including the prediction and gradient of the model. Recently, most white-box attack works adopt point-based attacks [34, 64, 53, 49]. Liu et al. [34] extended the gradient-based adversarial attack FGSM [17] strategy to point clouds, iteratively searching the desired point-wise perturbation under an $l_{2}$ norm constraint. 3D-Adv [58] introduced the C&W attack framework [7] in point cloud attacks, producing adversarial examples by shifting point coordinates and adding additional points with the guidance of the target classifier. Afterward, Tsai et al. [53] improved the C&W attack framework by introducing a KNN regularization term to suppress outlier points and compact point cloud surface. GeoA3 [55] uses a combined geometry-aware objective to maintain local curvature consistency and a uniform surface on the adversarial point cloud. Zheng et al. [65] proposed that deleting a small number of points with high saliency can effectively cause misclassification.
Apart from the point-based methods mentioned above, several studies have explored shape-based attack. Liu et al. [35] introduced shape-based attacks by adding new features to objects, as well as using a Gaussian kernel for sinking operations to deform point clouds. Zhang et al. [61] and Miao et al. [13] proposed directly attacking the mesh to generate smoother results, employing edge length regularization and Gaussian curvature regularization, respectively. Tang et al. [51] proposed to adversarially stretch the latent variables in an auto-encoder, which can be decoded as smooth adversarial point clouds. Lou et al. [36] proposed a shape-based attack HiT-ADV, which conducts a two-stage search for attack regions based on saliency and imperceptibility scores, and then adds deformation perturbations in each attack region using Gaussian kernel functions. Besides, some works [23, 46, 13, 51] attack point clouds in the feature space for imperceptible attack. While these methods eliminate outlier points and ensure smoothness, they still require optimization for each input point cloud, resulting in a high time cost for adversarial sample generation. To this end, universal attack [45, 12] was proposed to compute universal perturbations for point clouds with specific patterns.

### v3

@@ Abstract
Adversarial attacks exploit the vulnerability of deep models against adversarial samples. Existing point cloud attackers are tailored to specific models, iteratively optimizing perturbations based on gradients in either a white-box or black-box setting. Despite their promising attack performance, they often struggle to produce transferable adversarial samples due to overfitting to the specific parameters of surrogate models. To overcome this issue, we shift our focus to the data distribution itself and introduce a novel approach named NoPain, which employs optimal transport (OT) to identify the inherent singular boundaries of the data manifold for cross-network point cloud attacks. Specifically, we first calculate the OT mapping from noise to the target feature space, then identify singular boundaries by locating non-differentiable positions. Finally, we sample along singular boundaries to generate adversarial point clouds. Once the singular boundaries are determined, NoPain can efficiently produce adversarial samples without the need of iterative updates or guidance from the surrogate classifiers. Extensive experiments demonstrate that the proposed end-to-end method outperforms baseline approaches in terms of both transferability and efficiency, while also maintaining notable advantages even against defense strategies. Code and model are available at https://github.com/cognaclee/nopain.
@@ 2 Related Work
White-box attack on point cloud. Existing works of point cloud attacks can be roughly divided into white-box attacks and black-box attacks. Recently, most white-box attack works adopt point-based attacks [34, 64, 53, 49]. Liu et al. [34] extended the gradient-based adversarial attack FGSM [17] strategy to point clouds.3D-Adv [58] introduced the C&W attack framework [7] in point cloud attacks, producing adversarial examples by shifting point coordinates and adding additional points. Tsai et al. [53] improved the C&W attack framework by introducing a KNN regularization term to suppress outlier points and compact point cloud surface. GeoA3 [55] uses a combined geometry-aware objective to maintain local curvature consistency and a uniform surface.Zheng et al. [65] proposed that deleting a small number of points with high saliency can effectively cause misclassification. Apart from the point-based methods mentioned above, several studies have explored shape-based attack. Liu et al. [35] introduced shape-based attacks by adding new features to objects. Zhang et al. [61] and Miao et al. [13] proposed directly attacking the mesh to generate smoother results.Tang et al. [51] proposed to adversarially stretch the latent variables in an auto-encoder, which can be decoded as smooth adversarial point clouds. HiT-ADV [36] is a shape-based attack, that first search attack regions based on saliency and imperceptibility scores, and then adds deformation perturbations in each attack region with Gaussian kernel. Besides, some works [23, 46, 13, 51] attack point clouds in the feature space for imperceptible attack. While these methods eliminate outliers and ensure smoothness, they still require optimization for each point cloud, resulting in a high time cost. To this end, universal attack [45, 12] was proposed to compute universal perturbations for point clouds with specific patterns.
@@ 5 Conclusion
Acknowledgment
This research was supported by the National Key R&\&&D Program of China under Grant No. 2021YFA1003003, the Natural Science Foundation of China under Grant No. 62306059 and No. T2225012.